# On the degrees of freedom count on singular phase space submanifolds

### v3

@@ 3.1 The weak field limit of pure $R^{2}$ gravity
In this case, the equation of motion (25) gets reduced to [^2] Note that the corresponding result of Ref. [1] is different by the relative sign of the two terms in the formula (27). This is due to the opposite signature convention which reverts the sign in the definition of the $\square$ operator. I am grateful to the Authors of Ref. [1] for pointing this out to me, for initially I wrongly thought that there was some mistake of theirs.
@@
Of course, upon substitution of the solution for $\psi+\phi$ into the action, the latter turns out to be identically zero [1], precisely because the equation is $R=0$ while the Lagrangian is $R^{2}$. However, I would not use it as a proof that there is nothing else in the model. Substitution of a partial solution directly into the action might radically change its properties. The real proof is that, in the perturbative treatment, all the equations got reduced to only this one (27).
If to take the model of
@@
$\bigtriangleup\left(\phi+\psi\right)=-3\square\psi+b+c_{0}t+c_{i}x^{i}$ (29)
with constant numbers $b$ and $c_{\mu}$, which means a remaining full freedom of a spatially harmonic function with arbitrary dependence on time, and also one and there halves global degrees of freedom. Or more precisely, without relying on cosmological perturbations’ parametrisation, the $\bigtriangleup h_{00}$ quantity is given in terms of other components, with an addition of $b+c_{0}t+c_{i}x^{i}$.
On the other hand, if we remember that the linearised model is just an approximation, the unpredictability of a variable is not necessarily a gauge freedom. This is what also happens here. If we take a non-trivial spatially flat vacuum cosmology, then $R\neq 0$ and the extra scalar field takes the form of an ideal fluid in GR, therefore requiring $\phi=\psi$. If the limit of Minkowski space is taken while keeping this condition, then we have got a scalar degree of freedom. In other words, the count of degrees of freedom has a serious interpretational aspect here. There is no vanishing of the corresponding eigenvalue of the kinetic matrix. We simply have lost a constraint which used to make the initial data physical, in the meaning of influencing predictable outcomes.
Note also that, in what was called a manifestly covariant or à la Stückelberg approach, the paper [1] represented the metric perturbation in the form which can be taken roughly as
@@
It is not the standard Stückelberg procedure which introduces the new variables without restricting the old ones. The effect of this, only superficially ”à la Stückelberg”, approach is in putting extra derivatives on some parts of the metric, right into the action. The standard variational principle then deals with a weaker vanishing requirement, with more boundary conditions imposed on admissible variations, resulting in more general equations [5, 7]. I have also illustrated it on simple examples (11) in the previous Section. This is the real reason of the mismatch between different counts of degrees of freedom [1]. The model had got objectively modified by the parametrisation (30) put inside the action.
3.2 The 3-form toy model
Let me also use the 3-form model [1] for illustration of how their ”manifestly covariant” approach does change the model. In terms of the dual vector it has got an action
@@ 3.2 The 3-form toy model
In other words, the whole dynamics (32) means that
@@
The spatially transverse $A^{\mathcal{T}}_{i}$ part, $\partial_{i}A^{\mathcal{T}}_{i}\equiv 0$, represents the obvious gauge freedom. Otherwise, with certain interpretational issues again, we can say that the temporal component is a pure gauge one, too, while the longitudinal mode is then constrained by the equation (33). Therefore, no dynamical degrees of freedom, except one half global mode.
@@
in terms of the scalar field only, it is radically different, for the integrations by parts in terms of $\delta\phi$ require more boundary conditions than it was in terms of a 4D-longitudinal mode of $A_{\mu}$ itself, and therefore the extremality of the action is required for a limited class of variations only, producing equations of more freedom [5, 7]. The variation yields
@@
The Authors of the paper [1] correctly guessed that the problem lies in the partial ”gauge choice” when they put ${A^{T}}^{\mu}$ instead of just $A^{\mu}$ into the Stückelberg trick. This is a general fact. Fixing a gauge directly inside the action generically changes the theory. It wouldn’t have happened if we had done it already in the equations of motion. However, the variational prescription gets changed. In this particular case, the variation $\delta A_{\mu}$ vanishing at infinity is not equivalent to $\delta A^{T}_{\mu}$ and $\delta\phi$ vanishing there when $A_{\mu}=A^{T}_{\mu}+\partial_{\mu}\phi$.
On the other hand, it also sounds as a very strange aim, to ”obtain an action that is given just in terms of the gauge-invariant variables” [1]. How could one ever do that even for electrodynamics, let alone gravity? The gauge symmetry is a very important ingredient for writing an action. The equations can be written in terms of the field strength $F_{\mu\nu}$ only, however not the action itself since the action of $F_{\mu\nu}F^{\mu\nu}$ for $F_{\mu\nu}$ as a basic variable without charges would result in $F_{\mu\nu}=0$ only, with no electromagnetic waves [^3] This is because the gauge symmetry contains derivatives. Putting more derivatives into the action functional, like in a gauge-fixed Stückelberg trick, makes equations weaker and allows for more modes. Removing derivatives is opposite, it trivialises equations by making them stronger. A totally different example though is the coincident gauge of covariant teleparallel models [14].. One could of course write an action for the spatially transverse components only since those are invariant and obey a simple wave equation, but that would be a bit of cheating since it is then a different theory which totally ignores the Coulomb field and the Gauß law.
@@ 4 How (not) to perform the Hamiltonian analysis
In the paper [1], they also confirmed their findings by using the Hamiltonian analysis. With the correct result on the number of dynamical degrees of freedom, unfortunately this analysis also modified the contents of the models, for they went for extended Hamiltonians. I would like to use this occasion to explain, see also my paper [15], that the Dirac’s conjecture is always wrong, even in the cases when he thought it was correct, and therefore the secondary first class constraints should not be added to the Hamiltonian, unless we accurately redefine our model.
The main point is that, in most interesting cases, the first class constraints are not generators of gauge transformations. Only special combinations of the primary and the secondary ones are [16]. At the same time, adding the secondary constraints to the Hamiltonian increases the number of arbitrary Lagrange multipliers beyond the actual amount of gauge symmetry. And this is wrong, unless we admit that the momenta lose their direct relation with the fields and grab the physical meaning of the constrained variables which then become pure gauge.
In most expositions, this is simply swept under the rug as if only dynamical modes were physical. In electrodynamics, it then means [15] that the temporal and spatially longitudinal components of the vector-potential can be gauge transformed independently of each other, thus stripping the Coulomb potentials of physical significance. In gravity, the same injustice happens [15] to the scalar perturbations of the metric which are actually observed in temperature fluctuations of CMB.
@@ 4.1 The Hamiltonian of electrodynamics
In a sense, it is similar to a Stückelberg trick. We artificially enhance the space of variables by pronouncing the momenta independent. After that we add yet another gauge symmetry by simply calling anything unphysical as long as it does not commute with any of the first-class constraints. This is a well-defined symmetry. The constraints act independently in the sense of commuting with each other. Unlike second-class ones, they cannot fully restrict a coordinate and its own momentum simultaneously. In the example of electrodynamics, only momenta are in the constraints, therefore the corresponding transformations act only on the fields not touching the momenta which can then serve as new gauge-invariant quantities being equal to their initial non-invariant definition in a particular gauge.
The same happens in general. If the constraints are of the first class, then in each canonical pair of variables, they can involve only one combination and then transform another one. Therefore, for each canonical pair there must exist at least one gauge-invariant combination which can then serve as a substitute for the corresponding field. And by the very definition of the new gauge symmetry, the dynamics of invariant variables are not changed by adding the secondary constraints to the Hamiltonian. Therefore, in other cases of regular structure, it should also work [^4] See however the Ref. [17].. If we believe that the Hamiltonian mechanics is no less fundamental than the Lagrangian one, there is nothing wrong in considering such systems, too. I have nothing against this formal procedure as long as we honestly admit that we drastically change a model in order to get this result.
@@
Unfortunately, an extended Hamiltonian is often claimed to be the only appropriate one for gauge theories, without admitting the change in the dynamical system and without explaining that important parts of physical information got transported to momenta. It often leads to rather strange ideas. Sometimes people say that only two modes of electromagnetic field are physical, ignoring the fact there is a third physical, even though constrained, mode in the Coulomb potential. The same is for gravity. In GR, there are two dynamical modes and four more physical ones [15]. If we go for an extended Hamiltonian, it is very important to remember that these constrained physical modes do not disappear but rather get rewritten in new terms, and then not only the fields but also momenta must be taken as independent and potentially physical variables.
4.2 On the Proca fields with variable mass
Coming back to the models with ill-defined numbers of degrees of freedom, the singular locus of the variable-mass Proca field (12) is precisely $\psi=0$. The two constraints which always exist in this model, $\pi_{0}=0$ and $\partial_{i}\pi_{i}+2\psi A_{0}=0$ get their Poisson bracket vanish there and only there. When $\psi<0$, the model has got a ghost. The traditional way of seeing it is via a Stückelberg trick of $A_{\mu}\longrightarrow A_{\mu}+\partial_{\mu}\phi$ which makes the longitudinal mode a full-fledged ghost, for $\phi$ is then given a negative kinetic energy. If to not trust these tricks with derivative changes of variables, one can solve for the pair of the second-class constraints.
I am not interested in the ghosts right now, as long as they are stably there in the spectrum. Therefore, let me only briefly remind the reader about the ghost of the tachyonic vector field. For simplicity, consider just a standard Proca field (4) with $m^{2}<0$. One easily finds the constraints, the primary $\pi_{0}=0$ and the secondary $A_{0}=-\frac{1}{m^{2}}\partial_{i}\pi_{i}$ ones, which can easily be solved for, and upon substitution also produce an unbounded kinetic part $\frac{1}{2}\left(\pi^{2}_{i}+\frac{1}{m^{2}}(\partial_{i}\pi_{i})^{2}\right)$ of the Hamiltonian. I would say, it is still an interesting question, how dangerous this instability is, for predictability of (non-linear) classical evolution. However, this is beyond the scope of this paper.
It is also interesting to look at the Hamiltonian picture of the non-linear Proca model (17) with the potential energy of $\frac{1}{4}(A_{\mu}A^{\mu})^{2}$. We find a canonical Hamiltonian
the primary constraint
$\Phi=\pi_{0},$ (37)
and the secondary one
Generically, those (37,38) are the only pair of second-class constraints. However, we immediately see that their Poisson bracket
$\{\Phi,\tilde{\Phi}\}=3A^{2}_{0}-A^{2}_{i}$
@@ 4.3 The 3-form example
immediately reproduce the equation (33) and the definitions of momenta above. Of course, the secondary constraints are immediately obvious from the equation for $\dot{\pi}_{i}$. In this sense, they are already in the system.

### v5

@@ 3.1 The weak field limit of pure $R^{2}$ gravity
In this case, the equation of motion (25,27) gets reduced to [^3] Note that the corresponding result of Ref. [1] is different by the relative sign of the two terms in the formula (28). This is due to the opposite signature convention which reverts the sign in the definition of the $\square$ operator. I am grateful to the Authors of Ref. [1] for pointing this out to me, for initially I wrongly thought that there was some mistake of theirs.
@@
The paper [1] observes that, upon substitution of the solution for $\psi+\phi$ into the action, the latter turns out to be identically zero, in order to claim that there is nothing more in the physical spectrum. This fact is trivially true because the equation is $R=0$ while the Lagrangian is $R^{2}$. However, I would not use it as a proof that there is nothing else in the model. Substitution of a partial solution directly into the action might radically change its properties. The real proof is that, in the perturbative treatment, all the equations got reduced to only this one (28).
Perturbatively, one can uniquely find the field $\phi$ in terms of $\psi$. If to take the model of
@@
$\bigtriangleup\left(\phi+\psi\right)=-3\square\psi+b+c_{0}t+c_{i}x^{i}$
with constant numbers $b$ and $c_{\mu}$, which means a remaining full freedom of a spatially harmonic function with arbitrary dependence on time, and also one and three halves global degrees of freedom.
More precisely, without relying on cosmological perturbations’ parametrisation, the theory (29) has got an equation of motion
$\partial_{\alpha}\partial_{\beta}\left(\partial_{\mu}\partial_{\nu}-\eta_{\mu\nu}\square\right)h^{\mu\nu}=0$
which can be solved as $\left(\partial_{\mu}\partial_{\nu}-\eta_{\mu\nu}\square\right)h^{\mu\nu}=b+c_{0}t+c_{i}x^{i}$ and therefore implies that the $\bigtriangleup h_{00}$ quantity can be given in terms of other metric components, with an addition of $b+c_{0}t+c_{i}x^{i}$. Modulo the global addition and the freedom of a harmonic function, it can be taken as almost everything pure gauge and one constrained physical variable.
On the other hand, if we remember that the linearised model is just an approximation, the unpredictability of a variable is not necessarily a true gauge freedom. This is what happens here. However, while the tensor and vector perturbations have totally disappeared from the quadratic action, the story of the scalar mode is a more intricate one. If we take a non-trivial spatially flat vacuum cosmology, then $R\neq 0$ and the extra scalar field takes the form of an ideal fluid in GR, therefore requiring $\phi=\psi$. If the limit of Minkowski space is taken while keeping this condition, then we have got a scalar degree of freedom. In other words, the count of degrees of freedom has a serious interpretational aspect here. There is no vanishing of the corresponding eigenvalue of the kinetic matrix. We simply have lost a constraint which used to make the initial data physical, in the meaning of influencing predictable outcomes.
Note though that the locus of $R=0$ is non-perturbatively problematic. Indeed, any metric with $R\equiv 0$, i.e. subject to just one scalar equation, solves the full non-linear system of equations (22).
3.2 A parody of the Stückelberg trick
Recall also that the paper [1] reported a contradictory result, namely that they still saw a scalar mode in the pure $R^{2}$ theory linearised around Minkowski in what was called a manifestly covariant or à la Stückelberg approach. In all other cases of using this approach, the problem was simply in fixing a gauge right inside the action. Now it is slightly more problematic. They represented the metric perturbation in the form which can be taken roughly as (more precisely, see below)
@@ 3.2 A parody of the Stückelberg trick
As has been mentioned before, even if there were some gauge symmetry considerations behind, it would definitely not be the standard Stückelberg procedure which introduces new variables without restricting the old ones (by zero trace or by zero divergence). However, in this case, it is even worse. With the classical Stückelberg trick, the gauge-invariant combination is immediately obvious, while ambiguity of the spin formalism is much less transparent.
We can immediately calculate the (linearised) Ricci scalar (24) to be $R=-3\square\Psi$. If, in perturbation theory again, we treat the equation (25) as $R=0$, we get a dynamical mode of $\square\Psi=0$. The reason is that this representation (30), taken [1] from a previous paper [18], is very ambiguous [^4] One might worry that this parametrisation looks very similar to cosmological perturbation theory. However, the latter works in terms of spatial derivatives only, and ambiguities of solving equations for Laplacians are taken care of by boundary conditions at spatial infinity which is absolutely reasonable when in perturbation theory.. Let’s, for example, take the metric (26) which did not lead to any dynamical mode and find the trace and the double divergence of its perturbation for presenting it in the shape of (30). We get
$\square\Sigma+4\Psi=2\phi-6\psi\qquad\mathrm{and}\qquad\square^{2}\Sigma+\square\Psi=2\ddot{\phi}+2\bigtriangleup\psi,$
and can express $-3\square\Psi$ in terms of $\phi$ and $\psi$. In other words, the two Cauchy data for $\Psi$ simply correspond to a freedom of choosing a presentation of the metric in the shape of the formula (30).
At the same time, the action (29) turns into $9\int d^{4}x\cdot\left(\square\Psi\right)^{2}$ and produces even a 4-th order equation of motion $\square^{2}\Psi=0$. In this case, the field $\Psi$ was not given any extra derivatives in the action, however the operator of $\partial_{\mu}\partial_{\nu}$ had effectively been removed from it producing $\square R=0$ instead of $\partial_{\mu}\partial_{\nu}R=0$. Therefore, we get more solutions than the initial model used to have. It is yet another example of problems with substitutions right into the action.
Finally, to be precise, the scalar part of the metric perturbation (30) was taken [1, 18] in a bit different form of
$\left(\partial_{\mu}\partial_{\nu}-\frac{1}{4}\eta_{\mu\nu}\square\right)\mu+\frac{1}{4}\eta_{\mu\nu}\lambda=\partial_{\mu}\partial_{\nu}\mu+\frac{1}{4}\eta_{\mu\nu}\left(\lambda-\square\mu\right),$
I guess, with motivations from spin projector formalism so that the derivative part be traceless, which has of course given them a Lagrangian density proportional to $\left(\lambda-\square\mu\right)^{2}$, or $\left(\square\Psi\right)^{2}$ for the physical variable $\Psi=\lambda-\square\mu$. Being of a higher derivative order, such an action corresponds to two dynamical degrees of freedom, precisely like we have discussed above. They claimed [1] that one of those can be removed by the remnant gauge freedom of the Stückelberg trick, whatever it be in this case. It is certainly true in the sense of ambiguity of the chosen representation for the metric. However, what to make out of the second mode here? There are only very vague words, about the remnant gauge freedom again, in that paper [1] about it. Let me stress it once more that what has happened in reality is an adverse effect of making a substitution into the action functional.
3.3 A 3-form toy model
Let me also use the 3-form model from the same paper [1] for illustration of how their ”manifestly covariant” approach does change the model. In terms of the dual vector, the model has got an action
@@ 3.3 A 3-form toy model
In other words, the whole dynamics (32) of the theory (31) means that
@@
The spatially transverse $A^{\mathcal{T}}_{i}$ part, $\partial_{i}A^{\mathcal{T}}_{i}\equiv 0$, represents the obvious gauge freedom. Otherwise, we can say that the temporal component is a pure gauge one, too, while the longitudinal mode is then constrained by the equation (33). Therefore, no dynamical degrees of freedom, except one half global mode.
@@
in terms of the scalar field only, it is radically different, for the integrations by parts in terms of $\delta\phi$ need more boundary conditions than it was in terms of a 4D-longitudinal mode of $A_{\mu}$ itself, and therefore the extremality of the action is required for a limited class of variations only, producing equations of more freedom [6, 9]. The variation yields
@@
The Authors of the paper [1] correctly guessed that the problem lies in the partial gauge choice when they put ${A^{T}}^{\mu}$ instead of just $A^{\mu}$ into the Stückelberg trick. However, they describe one mode coming from the remnant freedom without any proper explanation for another mode residing inside the 4th order equation. As we have seen above, this is due to incorrectness of fixing a gauge. This is a general fact. Fixing a gauge directly inside the action generically changes the theory. It wouldn’t have happened if we had done it already in the equations of motion. However, the variational prescription gets changed. In this particular case, the variation $\delta A_{\mu}$ vanishing at infinity is not equivalent to $\delta A^{T}_{\mu}$ and $\delta\phi$ vanishing there when $A_{\mu}=A^{T}_{\mu}+\partial_{\mu}\phi$.
On one hand, fixing a gauge directly inside the action might not be a good idea. On the other hand, it also sounds as a very strange aim, to ”obtain an action that is given just in terms of the gauge-invariant variables” [1]. How could one ever do that even for electrodynamics, let alone gravity? The gauge symmetry is a very important ingredient for writing an action. The equations can be written in terms of only the field strength $F_{\mu\nu}$, however not the action itself since the action of $F_{\mu\nu}F^{\mu\nu}$ for $F_{\mu\nu}$ as a basic variable without charges would result in $F_{\mu\nu}=0$ only, with no electromagnetic waves [^5] This is because the gauge symmetry contains derivatives. Putting more derivatives into the action functional, like in a gauge-fixed Stückelberg trick, makes equations weaker and allows for more modes. Removing derivatives is opposite, it trivialises equations by making them stronger. A totally different example though is the coincident gauge of covariant teleparallel models [19].. One could of course write an action for the spatially transverse components only since those are invariant and obey a simple wave equation, but that would be a bit of cheating since it is then a different theory which totally ignores the Coulomb field and the Gauß law.
@@ 4 How (not) to perform the Hamiltonian analysis
In the paper [1], they also confirmed their findings by using the Hamiltonian analysis. With the correct result on the number of dynamical degrees of freedom, unfortunately this analysis also modified the contents of the models, for they went for extended Hamiltonians. I would like to use this occasion to explain, see also my paper [20], that unless we accurately redefine a model with secondary first-class constraints, the Dirac’s conjecture is always wrong, even in cases when it is commonly taken as correct, and therefore the secondary first class constraints should not be added to the Hamiltonian,
The main point is that, in most interesting cases, the first class constraints are not generators of gauge transformations. Only special combinations of the primary and the secondary ones are [21]. At the same time, adding the secondary constraints to the Hamiltonian increases the number of arbitrary Lagrange multipliers beyond the actual amount of gauge symmetry. And this is wrong, unless we admit that some momenta lose their direct relation with the fields and grab the physical meaning of the constrained variables which then become pure gauge.
In most expositions, this is simply swept under the rug as if only dynamical modes were physical. In electrodynamics, it then means [20] that the temporal and spatially longitudinal components of the vector-potential can be gauge transformed independently of each other, thus stripping the Coulomb potentials of physical significance. In gravity, the same injustice happens [20] to the scalar (and vector) perturbations of the metric which are actually observed in temperature fluctuations of CMB.
@@ 4.1 The Hamiltonian of electrodynamics
4.2 The new life of momenta
In a sense, what we saw about extending the Hamiltonian of electrodynamics is similar to the Stückelberg trick. We artificially enhance the space of variables by pronouncing the momenta independent. After that we add yet another gauge symmetry by simply calling anything unphysical as long as it does not commute with any of the first-class constraints. This is a well-defined symmetry. The constraints act independently in the sense of commuting with each other. Unlike second-class ones, they cannot fully restrict a coordinate and its own momentum simultaneously. In the example of electrodynamics, only momenta are in the constraints, therefore the corresponding transformations act only on the fields not touching the momenta which can then serve as new gauge-invariant quantities being equal to their initial non-invariant definition in a particular gauge.
The same happens in general. If the constraints are of the first class, then in each canonical pair of variables they can involve only one combination and transform another one. Therefore, for each canonical pair there must exist at least one gauge-invariant combination which can then serve as a substitute for the corresponding field. And by the very definition of the new gauge symmetry, the dynamics of invariant variables are not changed by adding the secondary constraints to the Hamiltonian. Therefore, in other cases of regular structure, it should also work [^6] See however the Ref. [22].. If we believe that the Hamiltonian mechanics is no less fundamental than the Lagrangian one, there is nothing wrong in considering such systems, too. I have nothing against this formal procedure as long as we honestly admit that we drastically change a model in order to get this result.
@@ 4.2 The new life of momenta
Unfortunately, an extended Hamiltonian is often claimed to be the only appropriate one for gauge theories, without admitting the change in the dynamical system and without explaining that important parts of physical information got transported to momenta. It often leads to rather strange ideas. Sometimes people say that only two modes of electromagnetic field are physical, ignoring the fact there is a third physical, even though constrained, mode in the Coulomb potential. It is important to remember that, in the extended Hamiltonian picture, it got moved to the sector of momenta.
The same is for gravity. In GR, there are two dynamical modes and four more physical ones [20]. If we go for an extended Hamiltonian, it is very important to remember that these constrained physical modes do not disappear but rather get rewritten in new terms, and then not only the fields but also momenta must be taken as independent and potentially physical variables. The primary constraints (the momenta of lapse and shift) are not generators of diffeomorphisms, for the coordinate changes are not about changing these functions alone. There are eight first-class constraints and only four arbitrary coordinates.
Still, I have also heard from some colleagues that there are only two physical components of the metric. In the Lagrangian, or total Hamiltonian formalism, this is certainly not true. The four constrained physical variables have got a different representation in the extended Hamiltonian description, though. However, it is very important to remember that they are there, especially if we are going to generalise the model. As one of the most trivial examples ever, we may generalise vacuum GR by adding matter contents, or just test particles. For the latter, the usual way is to look at the geodesics. And in this language, it is very important to have the full metric, in all its six physical components, and not just the two polarisations of the gravitational waves.
To sum it up, it’s not that one can prove that all first class constraints correspond to gauge symmetries. We can reformulate the model, though, by artificially extending the gauge symmetry so that they all indeed generate the new gauge transformations. An important point to bear in mind is that the physical information on the constrained modes gets then transferred from the fields to (their combinations with) momenta. When studying problematic loci of the phase space or generalising a model, one should not forget about the differences.
@@ 4.3 The 3-form example
immediately reproduce the equation (33) and the definitions of momenta above. Of course, the secondary constraints are obvious from the equation for $\dot{\pi}_{i}$. In this sense, they are already in the system.
@@
4.4 On the Proca fields with variable mass
Coming back to the models with ill-defined numbers of degrees of freedom, the singular locus of the variable-mass Proca field (12) is precisely $\psi=0$. The two constraints which always exist in this model, $\pi_{0}=0$ and $\partial_{i}\pi_{i}+2\psi A_{0}=0$ get their Poisson bracket vanish there and only there. When $\psi<0$, the model has got a ghost. The traditional way of seeing it is via a Stückelberg trick of $A_{\mu}\longrightarrow A_{\mu}+\partial_{\mu}\phi$ which makes the longitudinal mode a full-fledged ghost, for $\phi$ is then given a negative kinetic energy, or one can solve for the pair of the second-class constraints.
I am not interested in the ghosts right now, as long as they are stably there in the spectrum. Therefore, let me only briefly remind the reader about the ghost of the tachyonic vector field. For simplicity, consider just a standard Proca field (4) with $m^{2}<0$. One easily finds two constraints, the primary $\pi_{0}=0$ and the secondary $A_{0}=-\frac{1}{m^{2}}\partial_{i}\pi_{i}$ ones, which can easily be solved for, and upon substitution also produce an unbounded kinetic part $\frac{1}{2}\left(\pi^{2}_{i}+\frac{1}{m^{2}}(\partial_{i}\pi_{i})^{2}\right)$ of the Hamiltonian. I would say, it is still an interesting question, how dangerous this instability is, for predictability of (non-linear) classical evolution. However, this is beyond the scope of this paper.
It is also interesting to look at the Hamiltonian picture of the non-linear Proca model (17) with the potential energy of $\frac{1}{4}(A_{\mu}A^{\mu})^{2}$. We find the canonical Hamiltonian
the primary constraint
$\Phi=\pi_{0},$ (41)
and the secondary one
Generically, those (41,42) are the only pair of second-class constraints. However, we immediately see that their Poisson bracket
$\{\Phi,\tilde{\Phi}\}=3A^{2}_{0}-A^{2}_{i}$
What we see here was dubbed ”constraint bifurcation” in the context of metric-affine generalisations of gravity [24, 25]. These troubles are indeed very often much easier detected in the Hamiltonian approach. However, it is not then that clear how to deal with them. In the Lagrangian formalism, we do at least have a partial differential equation whose ill-posed Cauchy problem can be analysed. In Hamiltonian mechanics, the theory of non-degenerate Lagrangian systems is well-known. It is beautiful and mathematically rigorous. However, for the constrained systems, it is mostly presented rather as a recipe, which actually assumes constant ranks of the Poisson brackets’ algebras of constraints. It is not clear, how to use it for calculating a number when the number is not well-defined, to start with.